• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name R. Michaels

    Log entry time 00:52:22 on June13,2010

    Entry number 326149

    keyword=rates on BeO unstable

    Restore raster to 23 x 9.5 (MCC units).
    Move to BeO.
    Restoer laser = 65, slit = 15. Here we had 13 kHz before

    Rates on BeO are wildly unstable. For both HRS.
    Investigating.

    At least its not the septum.

    -----------

    After a time, MCC decides it may be a steering problem.
    Turned off HV. Get 10 uA on BeO. They will steer, etc.


    FIGURE 1